OpenStudy (anonymous):

please help me guys = ]!!! Given two spheres, if the volume of the first sphere is 4/3*3.14 and the volume of the second sphere is 36*3.14 , which is the relation of the second radius to the first? squared tripled doubled multiplied by a factor of 8/3

OpenStudy (anonymous):

First lets find out how many times the volume grew. We will divide the second sphere's volume by the first sphere: \[\frac{36*\pi}{\frac{4}{3}*\pi}\] This comes out to be 27 (i think, might want to plug that in a calculator.) So the volume grew by a factor of 27, but thats in cubic units (volume is measured in cubic units). A radius is just regular units, so we cube root 27 to get out real answer: \[\sqrt[3]{27} = \sqrt[3]{3^3} = 3\] So tripled is the answer.

OpenStudy (anonymous):

thanks hun you helped me when no one could!